Use the convolution theorem to find the inverse Laplace transform of the given function. 6 ਚ s3 (2) +0 (t) = (s +9)
Answer #1

F(s)=\frac{6}{s^3(s^2+9)}

F(s)=(\frac{6}{s^3})(\frac{1}{s^2+9})=g(s)h(s)

Where

g(s)=\frac{6}{s^3}

h(s)=\frac{1}{s^2+9}

We know

L^{-1}(F(s))=L^{-1}(g(s)h(s))=L^{-1}(g(s))*L^{-1}(h(s)). (1)

Where * represents convolution.

L^{-1}(g(s))=L^{-1}(\frac{6}{s^3})=3t^2

And

L^{-1}(h(s))=L^{-1}(\frac{1}{s^2+9})=\frac{sin(3t)}{3}

By (1)

L^{-1}(F(s))=\int_{0}^{t} u^2 sin3(t-u) du=\frac{2cos(3t)+9t^2-2}{27}

so

L^{-1}(F(s))=\frac{2cos(3t)+9t^2-2}{27}

answer.